OpenStudy (anonymous):

Solve five-twelfths m plus two-thirds equals negative three-fourths for m. m equals negative one-fifth m equals negative five over one hundred forty-four m equals negative three and two-fifths m equals negative eighty-five over one hundred forty-four

OpenStudy (goldphenoix):

\[\frac{ 5 }{ 12 }m + \frac{ 2 }{ 3 } = -\frac{ 3 }{ 4 }\] You want to find m, so you want to get rid of 2/3. So do: \[\frac{ 3 }{ 4 } - \frac{ 2 }{ 3 } = \frac{ 1 }{ 12 }\] Now you will have: \[\frac{ 5 }{ 12 }m = \frac{ 1 }{ 12 }\] The final step to solving for m is to divide 5/12 from 1/12. \[\frac{ 1 }{ 12 } \div \frac{ 5 }{ 12 } = \frac{ 1 }{ 5 }\]

OpenStudy (goldphenoix):

Forgot to add the negative for -3/4. So it would be - 1/5, I believe.

OpenStudy (anonymous):

Thank you so much!

OpenStudy (goldphenoix):

I hope this help. :)

OpenStudy (goldphenoix):

Hm... Wait. Let me check something, I think the negative mest it up.